OpenStudy (anonymous):

Suppose f( x)= x^2 + x +3 Compute the following: a. f ( -2)+ f(4) b. f ( -2) - (4)

OpenStudy (anonymous):

f(-2) = (-2)^2 +(-2) + 3 = 4 - 2 + 3 = 5 f(4) = 4^2 +4 +3 = 16 +4 +3 = 23 a) 28 b) -18